Find the x-intercepts
Y= x^2+4x-1

Answer by actuary(112)

The x-intercepts of the equation can be found by finding the roots of the equation x^2+4x-1. Use the quadratic formula x =( -b +/- Square Root[b^2-4ac])/(2*a)
where a= 1, b=4 and c=-1.
Sunstituting these values intp the quadratic formula we have
-4 +/- Square Root[16-4*1*(-1) = =4 +/- Square Root[20]
You can simplifly Square Root[20]=Square Root[4]*Square Root[[5]=Square[2^2]*Square[5]=2*Square Root[5]
The the x-intercepts are (-4+2*Square[5])/2 = -2+Square Root[5] AND
-2-Square Root[5].
In both cases, I was able to cancel 2 in the ny=umerator and denominator.
Learning how to simplify the Sqyare Root of a number was probably one of the objectuves of this problem along with practice using the quadatic formula. Also, don't forget that a negative number times anegative number is a positive number.